OpenStudy (anonymous):

Turn the Riemann Sum (-1+k(3/n))^4 3/n into a definite integral

OpenStudy (anonymous):

I have the limits, I know they go from -1 to 2, but I'm not sure how to find f(x)

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty} \sum_{k=1}^{n}f(a+i \Delta x) \cdot \Delta x=\int\limits_{a}^{b}f(x) dx \text{ where } \Delta x=\frac{b-a}{n}\]

OpenStudy (freckles):

so you say a is -1 and bis 2 so delta x is (2-(-1))/n=3/n so you have \[\lim_{n \rightarrow \infty} \sum_{i=1}^{n}(a+i \Delta x)^4 \cdot \Delta x \] which is the exact form on the left hand side of the equality above you just have to identify f(x) now

OpenStudy (anonymous):

Maybe I wrote it wrong, but I believe that delta x is 3/n . So would f(x) be 3x^4?

OpenStudy (freckles):

yes delta x is 3/n

OpenStudy (freckles):

so I replaced the 3/n 's with delta x

OpenStudy (freckles):

you mean f(x)=x^4

OpenStudy (freckles):

Deltax is 3/n not 1/n

OpenStudy (anonymous):

Okay, so what is inside the parentheses is x in f(x)?